S. Mr. X is known to hit the target in 4 out of 5 shots, where as Mr. Y is known to hit the target in 3 out of 5 shots. Find the probability that both will hit the Target


1
Expert's answer
2022-01-18T10:48:51-0500

Let A be "both hit", H1 - "X hit", H2 - "Y hit", then, since X and Y shoots independently

"P(A)=P(H1)*P(H2)={\frac 4 5}*{\frac 3 5}={\frac {12} {25}}=0.48"